Jones Company had 100 units in beginning inventory at a total cost of $10,000.The company
purchased 200 units at a total cost of $26,000. At the end of the year, Jones had 80 units in ending inventory.

Compute the cost of the ending inventory and the cost of goods sold under (1) FIFO, (2)
LIFO, and (3) average-cost.
Answer #1
Since this is year-end, it will be assumed to be a periodic inventory system, so the exact date on which the goods were sold is not important.

(1) FIFO
Under FIFO (first in-first out), goods are assumed to be sold in order of purchase, so is the cost.
Purchases during the period:
100 units at $100 = $10,000
200 units at $130 = $26,000

Units sold during the period = 220
Cost of units sold
=100*$100+120*130=$25,600
Value of ending inventory
=10,000+26,000-25,600
=$10,400

(2) LIFO
Under LIFO (last in-first out), goods purchased last is sold first.
Purchases during the period:
100 units at $100 = $10,000
200 units at $130 = $26,000

Units sold during the period = 220
Cost of units sold
=20*$100+200*130=$28,000
Value of ending inventory
=10,000+26,000-28,000
=$8,000

(3) average-cost
Under average-cost, cost of goods is averaged over the period and the cost applies to the whole period.
Purchases during the period:
100 units at $100 = $10,000
200 units at $130 = $26,000
average cost per unit
=(10,000+26,000)/300
=$120 per unit

Units sold during the period = 220
Cost of units sold
=220 * $120
=$26,400
Value of ending inventory
=36,000-26,400
=$9,600
answered by: breeawna